OpenStudy (anonymous):

If f(x)=x^2+3x+5 what is f(3+h)?

OpenStudy (anonymous):

\[f(3+h)_{?}\]

OpenStudy (saifoo.khan):

\[f(x)=x^2+3x+5\] \[f(3+h)=(3+h)^2+3(3+h)+5\]

OpenStudy (anonymous):

that's not simplified right?

OpenStudy (anonymous):

is it (3+h)^2+8+h ???

OpenStudy (saifoo.khan):

no, that is not simplified.

OpenStudy (saifoo.khan):

(3+h)^2+3(3+h)+5 (3+h)^2 + 9 + 3h + 5 and so on..

OpenStudy (anonymous):

is my answer right?

OpenStudy (saifoo.khan):

no.

OpenStudy (saifoo.khan):

Blank, go to chat.

OpenStudy (anonymous):

never mind got it